OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 7)^2(x - 7)^3

OpenStudy (anonymous):

ok, so for what values of x does f(x)=0?

OpenStudy (anonymous):

I'll get you started. If x=-7 you have: f(x)=5(-7+7)^2(x-7)^3 f(x)=5(0)^2(x-7)^3 f(x)=0 Since the exponent on the (x+7) term is 2, we say that this zero has a "multiplicity of 2". You can find the other one I think now.

OpenStudy (anonymous):

note: I didn't bother inputting x=-7 to the second term above (it won't change anything anyways).

OpenStudy (anonymous):

how did you get (x+7) if the problem is \[f(x) = 5(x + 7)^{2}(x - 7)^{3}\] don't you have to solve (x+7)(x+7) and then (x-7)(x-7)(x-7) ?

OpenStudy (anonymous):

You don't have to expand this one. You just need to see what values of x will make the entire thing zero. If the (x+7)^2 term is zero, then the entire f(x) is zero because zero multiplied by anything is zero. So I noted that if x=-7, then (-7+7)^2=0. So, f(-7)=5(0)(-7-7)^3=0 So x=-7 is a zero. It has a multiplicity of two because it appears twice. f(x)=5(x+7)(x+7)(x-7)^3

OpenStudy (anonymous):

Now you have to look at the other term (x-7)^3 to find the other zero.